# Subgradient Convergence Implies Subdifferential Convergence on Weakly Convex Functions: With Uniform Rates Guarantees

### v1

@@ 5.1.1 Noiseless Phase Retrieval
For standard normal measurement vectors $a\sim\mathsf{N}(0,I)$, a recent important and interesting study first characterizes $\mathcal{Z}$, the locations of stationary points of the population objective $\Phi(x)$ [DDP20]. Remarkably, the authors further develop a quantitative version of Attouch’s epi-convergence theorem to show that the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt[4]{d/m}$DDP20]. More precisely, [DDP20, Theorem 5.2, Corollary 6.3] prove that, there exist numerical constants $c,C>0$ such that when $m\geq Cd$, with an inner probability at least $1-C\exp(-cd)$ (see Section 2.4 for a definition of inner probability):
@@
In words, with high probability, for any stationary point of $\Phi_{S}$, there is a stationary point of $\Phi$ that is at most a distance of $C\sqrt[4]{d/m}\left\|{\bar{x}}\right\|$ in the $\ell_{2}$ norm away from it. It is important to note that the slow rate in $m$, namely $1/\sqrt[4]{m}$, appears intrinsic to the approach based on the quantitative version of Attouch’s epi-convergence theorem [DD22, Section 5].
In this paper, we show a stronger result for large sample size $m$: the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt{d/m}$ up to logarithmic factors in $d,m$. This rate improvement in $m$ is due to our bound on uniform convergence of subdifferentials of $\Phi$ and $\Phi_{S}$ in Theorem 5.1, and our proof of Theorem 5.2 benefits from the characterization of approximate stationary points of the population objective $\Phi$ [DDP20, Corollary 5.3]. Nonetheless, our result, Theorem 5.2, comes with a tradeoff of additional logarithmic factors in $d$ and $m$.
Theorem 5.2.
@@
The formal proof of Theorem 5.2 is deferred to Appendix, Section B.
@@ 5.2 Illustration II: Matrix Sensing
Our second illustration demonstrates the flexibility of our framework. Consider the robust matrix sensing objective [LZMCSV20, CCD+21, DJC+21]:
@@
Another application of Theorem 4.2 yields an upper bound on the subdifferential gap for the robust matrix sensing objective. Recall a matrix $A\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian random matrix if the vectorized matrix $\textbf{{\rm vec}}(A)\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian vector in the sense of Definition 5.1 [Ver18].
Corollary 5.3.
@@
The proof is provided in Section C and is analogous to the phase retrieval case. The primary challenge, which comes from evaluating the VC dimension, is addressed using Theorem 6.1.
@@ 6 VC Dimension Bounds
The following theorem, rooted in the field of real algebraic geometry, is referenced from Matousek’s textbook on discrete geometry [Mat13]. See [Mil64, Tho65] and [War68, Theorem 3].
Theorem 6.2 ([Mat13, Theorem 6.2.1]).
@@
By definition, the number of possible binary patterns, $|\mathcal{B}|$, cannot exceed the number of sign patterns for the same polynomials $q_{1},q_{2},\ldots,q_{N}$. According to Theorem 6.2, the number of sign patterns is bounded by $(50KN/(d+1))^{d+1}$. Therefore, we conclude:
@@ 7 Discussions
This paper presents a general technique for proving subdifferential convergence in weakly convex functions through the establishment of subgradient convergence (Theorem 3.1). We demonstrate this approach to stochastic convex-composite minimizations—an important class of weakly convex objectives—and derive concrete convergence rates for subdifferentials of empirical objectives using tools from statistical learning theories (Theorem 4.2). Our results achieve optimal rates with tight dependence on the sample size and dimensionality (up to logarithmic factors) and do not rely on key distributional assumptions in the literature requiring the subdifferential mappings to be continuous under the Hausdorff metric. In Section 5, we demonstrate how our results lead to complementary understanding of subdifferential convergence and its uses for landscape analysis, in more specific applications such as robust phase retrieval and matrix sensing problems.
@@
  1. 1.
    Can Theorem 4.2 be extended to other weakly-convex models, such as max-of-smooth functions [Roc81, HL23]? Extending this theorem could provide tools for solving other weakly-convex optimization problems.
  2. 2.
    Can Theorem 3.1 be extended to other set-valued mappings, such as outer semicontinuous monotone operators? Extending this theorem could enhance our understanding of solutions to certain stochastic variational inequalities [FP03].
  3. 3.
    Theorem 4.2 could facilitate our understanding of the landscape of a certain class of nonsmooth empirical risk minimizations. Exploring its statistical applications, such as nonlinear quantile regressions [KP96, Koe05], may lead to new insights into data-driven models.
8 Acknowledgement
Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution.
Appendix A Proof of Theorem 4.1
In this section, we prove Theorem 4.1. The key to our proof is to establish high probability upper bounds of the following two quantities:
@@ 8 Acknowledgement
Given these high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2), Theorem 4.1 follows by plugging these high probability bounds into equation (A.1).
It remains to prove the high probability bounds on $\Gamma_{1}$ and $\Gamma_{2}$ as specified in equation (A.2). To unify the proofs for these two bounds, we utilize Theorem A.1, which details the uniform convergence rate involving indicator functions and smooth functions. The proof of Theorem A.1 is in Section A.1.
Theorem A.1.
@@
We will apply Theorem A.1 for two different sets of functions $o,w$ to deduce the high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2). Below we set $\mathcal{Z}_{1}=\mathbb{R}^{d}$ and $\mathcal{Z}_{2}=\mathbb{B}(x_{0};r)$.
First, we set $o(x,\xi)=c(x,\xi)$ and $w(x^{\prime},\xi)=\nabla c(x^{\prime},\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem A.1 are satisfied. Applying this theorem yields the following guarantee. Let
@@
Next, we set $o(x,\xi)\equiv 0$ and $w(x^{\prime},\xi)=(h^{sm})^{\prime}(c(x^{\prime};\xi))\cdot\nabla c(x^{\prime};\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem A.1 are satisfied. Applying this theorem provides the following guarantee. Let
@@
At this point, the proof of Theorem 4.1 is complete.
A.1 Proof of Theorem A.1
Our proof of Theorem A.1 combines two well-established techniques for proving uniform bounds on the supremum of a random process in the statistical learning theory: chaining, which constructs a sequence of progressively finer approximations to the continuous process, and Sauer-Shelah Lemma, a combinatorial principle using the VC dimension to control the complexity of the function class under consideration [vdVW96, Chapter 2]. The chaining argument specifically addresses the continuous parameter $z_{2}$, and the complexity of its increments is then controlled using the VC dimension technique applied to the function class specified by the parameter $z_{1}$. To deal with the measurability issues, we use the notion of outer integral [vdVW96, Chapter 1.2], and prove that our process is measurable after symmetrization, following the idea in [vdVW96, Chapter 2.3].
@@ A.1.1 Chaining Argument
Let $\epsilon_{k}=r2^{-k}$ for every $k\geq 0$. An $\epsilon_{k}$-cover of $\mathcal{Z}_{2}$, denoted by $\mathcal{Z}_{2,\epsilon_{k}}$, enables the approximation of every point $z_{2}\in\mathcal{Z}_{2}$ by a point $\pi_{k}(z_{2})\in\mathcal{Z}_{2,\epsilon_{k}}$ such that $\left\|{z_{2}-\pi_{k}(z_{2})}\right\|\leq\epsilon_{k}$. Specifically, we take the initial cover $\mathcal{Z}_{2,\epsilon_{0}}=\{z_{2,0}\}$ to consist solely of the center $z_{2,0}$ of the ball $\mathcal{Z}_{2}$. By employing a standard volume-type argument, we can further select the cover $\mathcal{Z}_{2,\epsilon_{k}}$ such that $\log|\mathcal{Z}_{2,\epsilon_{k}}|\leq p\log(2+r/\epsilon_{k})$ for every $k\geq 1$ [Ver18, Corollary 4.2.13].
Using these coverings, we can decompose $R_{m}(z_{1},z_{2})$ through a chaining argument [Dud67]:
@@
The proof of Lemma A.2 is deferred to Section A.1.2.
@@
for each $k\geq 1$. Below we use $C$ to denote the universal constant that appears in Lemma A.2.
First, by Lemma A.2, with an inner probability at least $21-\delta/$ 2:
@@
Next, the adjustment of $\delta$ into $\delta_{k}$ allows the use of a union bound to extend the probability guarantees from Lemma A.2 from individual pairs to all pairs within the cover sets $\mathcal{Z}_{2,\epsilon_{k}}$ and $\mathcal{Z}_{2,\epsilon_{k-1}}$. Specifically, for each $k\geq 1$, with an inner probability at least $1-\delta\cdot 2^{-(k+1)}$, there is the bound:
@@
holds with probability at least $1-\delta$. The proof of Theorem A.1 is then complete.
A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma A.2)
We give a unifying proof for the two statements in Lemma A.2. To simplify the notation, we define:
@@
We will first demonstrate how Lemma A.4 implies Lemma A.2 and defer the proof of Lemma A.4 to the end. Introduce two notation:
@@
Note the similarities between the expressions in equation (A.6) and in equation (A.7). Because $u_{z_{2},z_{2}^{\prime}}(\xi)$ and $u_{z_{2,0}}(\xi)$ are subexponential random vectors, the conditions of Lemma A.2 are met. More precisely, $u_{z_{2},z_{2}^{\prime}}(\xi)$ is $\sigma\left\|{z_{2}-z_{2}^{\prime}}\right\|$ subexponential for every pair of $z_{2},z_{2}^{\prime}\in\mathcal{Z}_{2}$, and $u_{z_{2,0}}(\xi)$ is $\sigma$ subexponential. This allows us to leverage Lemma A.4 to conclude that Lemma A.2 follows.
In the remainder, we prove Lemma A.4. Our proof exploits the fact that the VC dimension of the set family $\mathcal{W}$ controls the Rademacher complexity associated with the indicators $\psi_{z_{1},t}(\xi)$.
To deal with outer probability, we need the notion of outer integral [vdVW96, Chapter 1.2].
Definition A.1.
@@
Outer integrals exhibit monotone property: if $T_{1}(\xi)\leq T_{2}(\xi)$ for every $\xi$, then $\mathbb{E}^{*}[T_{1}]\leq\mathbb{E}^{*}[T_{2}]$. This is essentially all we need, as the remaining properties of outer integral will be reference from the textbook [vdVW96]. To ensure clarity when discussing the calculus rules of the outer integral, we will explicitly reference the relevant chapters in the textbook [vdVW96].
Proof of Lemma A.4.
We define the moment-generating function, using the notion of outer integral (see Definition A.1):
@@
Let $V_{1/2}$ denote a $1/2$ cover of unit ball in $\mathbb{R}^{d}$ with $\log|V_{1/2}|\leq d\log 6$. Following a standard argument (see, e.g., [Ver18, Exercise 4.2.2]), we know that for every $\bar{v}\in\mathbb{R}^{d}$:
@@
Furthermore, we introduce i.i.d. Rademacher random variables $\{\epsilon_{i}\}_{i=1}^{m}$ with $\mathbb{P}(\epsilon_{i}=\pm 1)=1/2$, independent of the data samples $\{\xi_{i}\}_{i=1}^{m}$. To simplify our notation, we use $\epsilon_{1:m}$ to denote the vector $(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{m})$ and $\xi_{1:m}$ to denote the vector $(\xi_{1},\xi_{2},\ldots,\xi_{m})$.
Applying symmetrization techniques from empirical process theory for the outer integral [vdVW96, Lemma 2.3.1], we deduce an upper bound for $U(\tau)$:
$\begin{split}U(\tau)\leq\mathbb{E}^{*}\left[\exp\left(4\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$ (A.10)
where $\mathbb{E}^{*}$ is understood as the outer integral on the product probability space involving the data samples $\xi_{1:m}$ and the independent Rademacher random variables $\epsilon_{1:m}$. To be more formal, we follow [vdVW96, Chapter 2.3]. We can think of our data samples $\xi_{1:m}$ being defined on the product probability space $(\Xi^{\otimes m},\sigma(\mathscr{G}^{\otimes m}),\mathbb{P}^{\otimes m})$, where $\Xi^{\otimes m}$ represents the $m$-fold Cartesian product of $\Xi$, $\sigma(\mathscr{G}^{\otimes m})$ represents the $\sigma$-field generated by the $m$-fold product of $\mathscr{G}$, and $\mathbb{P}^{\otimes m}$ is the $m$-fold product of $\mathbb{P}$, where $\xi_{1},\xi_{2},\ldots,\xi_{m}$ corresponds to the coordinate projections. Then, we introduce Rademacher random variables $\epsilon_{1:m}=(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{m})$ defined on another probability space $(\Xi^{\prime},\mathscr{H},\mathbb{P}^{\prime})$. We finally define the product probability space of interest:
@@
In this product probability space, the sample $\xi_{1:m}$ corresponds to the first $m$ coordinates projections and the additional Rademacher random variables $\epsilon_{1:m}$ depending only on the $(m+1)$st coordinate.
Lemma A.5 is the critical component that resolves the measurability issue. From the empirical process theory, we know that measurability of certain forms is needed at this point, as Fubini’s theorem is not valid for outer integrals [vdVW96, Chapter 2.3]. Lemma A.5 follows the established ideas to resolve the measurability issues, see, e.g., discussions in the textbook on the measurable class [vdVW96, Definition 2.3.3]. The proof of Lemma A.5 is deferred to Section A.1.4.
Lemma A.5.
@@
$(\xi_{1:m},\epsilon_{1:m})\mapsto\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})$
@@
Lemma A.5 provides the justification for replacing the outer integral $\mathbb{E}^{*}$ on the RHS of equation (A.10) with the standard integral $\mathbb{E}$. Indeed, equation (A.10) now leads to
$\begin{split}U(\tau)\leq\mathbb{E}\left[\exp\left(4\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$ (A.12)
@@
$\begin{split}V(\tau,\xi_{1:m})=\mathbb{E}_{\epsilon_{1:m}}\left[\exp\left(\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$
where the symbol $\mathbb{E}_{\epsilon_{1:m}}$ means the standard expectation taken solely over the Rademacher variables $\epsilon_{1:m}$. As the Fubini’s theorem holds for the standard expectation, under this notation, the inequality (A.12) then becomes:
@@
For a reference of Sauer-Shelah Lemma, see, e.g., [Sau72, She72] or [Ver18, Theorem 8.3.16].

### v2

@@ 5.1.1 Noiseless Phase Retrieval
For standard normal measurement vectors $a\sim\mathsf{N}(0,I)$, a recent important and interesting study first characterizes $\mathcal{Z}$, the locations of stationary points of the population objective $\Phi(x)$ [14]. Remarkably, the authors further develop a quantitative version of Attouch’s epi-convergence theorem to show that the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt[4]{d/m}$14]. More precisely, [14, Theorem 5.2, Corollary 6.3] prove that, there exist numerical constants $c,C>0$ such that when $m\geq Cd$, with an inner probability at least $1-C\exp(-cd)$ (recall the definition of inner probability from Section 2.4):
@@
In words, with high probability, for any stationary point of $\Phi_{S}$, there is a stationary point of $\Phi$ that is at most a distance of $C\sqrt[4]{d/m}\left\|{\bar{x}}\right\|$ in the $\ell_{2}$ norm away from it. It is important to note that the slow rate in $m$, namely $1/\sqrt[4]{m}$, appears intrinsic to the approach based on the quantitative version of Attouch’s epi-convergence theorem [16, Section 5].
In this paper, we show a stronger result for large sample size $m$: the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt{d/m}$ up to logarithmic factors in $d,m$. This rate improvement in $m$ is due to our bound on uniform convergence of subdifferentials of $\Phi$ and $\Phi_{S}$ in Theorem 6, and our proof of Theorem 7 benefits from the characterization of approximate stationary points of the population objective $\Phi$ [14, Corollary 5.3]. Nonetheless, our result, Theorem 7, comes with a tradeoff of additional logarithmic factors in $d$ and $m$.
Theorem 7.
@@
The formal proof of Theorem 7 is deferred to Appendix, Section B.
@@ 5.2 Illustration II: Matrix Sensing
Our second illustration demonstrates the flexibility of our framework. Consider the robust matrix sensing objective [10, 11, 41]:
@@
Another application of Theorem 5 yields an upper bound on the subdifferential gap for the robust matrix sensing objective. Recall a matrix $A\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian random matrix if the vectorized matrix $\textbf{{\rm vec}}(A)\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian vector in the sense of Definition 3 [44]. For clarity in notation, we use $\left\|{\cdot}\right\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{D\times D}$.
Corollary 1.
@@
The proof is provided in Section C and is analogous to the phase retrieval case. The primary challenge, which comes from evaluating the VC dimension, is addressed using Theorem 8.
@@ 6 VC Dimension Bounds
The following theorem, rooted in the field of real algebraic geometry, is referenced from Matousek’s monograph on discrete geometry [52]. See [50, 51] and [53, Theorem 3].
Theorem 9 ([52, Theorem 6.2.1]).
@@
By definition, the number of possible binary patterns, $|\mathcal{B}|$, cannot exceed the number of sign patterns for the same polynomials $q_{1},q_{2},\ldots,q_{N}$. According to Theorem 9, the number of sign patterns is bounded by $(50KN/(d+1))^{d+1}$. Therefore, we conclude:
@@ 7 Discussions
This paper presents a general technique for proving subdifferential convergence in weakly convex functions through the establishment of subgradient convergence (Theorem 1). We demonstrate this approach to stochastic convex-composite minimizations—an important class of weakly convex objectives—and derive concrete convergence rates for subdifferentials of empirical objectives using tools from statistical learning theories (Theorem 5). Our results achieve optimal rates with tight dependence on the sample size and dimensionality (up to logarithmic factors) and do not rely on key distributional assumptions in the literature requiring the subdifferential mappings to be continuous under the Hausdorff metric. In Section 5, we demonstrate how our results lead to complementary understanding of subdifferential convergence and its uses for landscape analysis, in more specific applications such as robust phase retrieval and matrix sensing problems.
@@
  • Can Theorem 5 be extended to other weakly-convex models, such as max-of-smooth functions [54, 55]? Extending this theorem could provide tools for solving other weakly-convex optimization problems.
  • Can Theorem 1 be extended to other set-valued mappings, such as maximal monotone operators? Extending this theorem could enhance our understanding of solutions to certain stochastic variational inequalities [56].
  • Theorem 5 could facilitate our understanding of the landscape of a certain class of nonsmooth empirical risk minimizations. Exploring its statistical applications, such as nonlinear quantile regressions [13, 57], may lead to new insights into data-driven models.
[…]
Acknowledgements Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution. He is also grateful to Damek Davis and Dmitriy Drusvyatskiy for their valuable feedback on the initial manuscript draft posted on arXiv.
Appendix A Proof of Theorem 4
In this section, we prove Theorem 4. The key to our proof is to establish high probability upper bounds of the following two quantities:
@@ Appendix A Proof of Theorem 4
Given these high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2), Theorem 4 follows by plugging these high probability bounds into equation (A.1).
It remains to prove the high probability bounds on $\Gamma_{1}$ and $\Gamma_{2}$ as specified in equation (A.2). To unify the proofs for these two bounds, we utilize Theorem 10, which details the uniform convergence rate involving indicator functions and smooth functions. The proof of Theorem 10 is in Section A.1.
Theorem 10.
@@
We will apply Theorem 10 for two different sets of functions $o,w$ to deduce the high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2). Below we set $\mathcal{Z}_{1}=\mathbb{R}^{d}$ and $\mathcal{Z}_{2}=\mathbb{B}(x_{0};r)$.
First, we set $o(x,\xi)=c(x,\xi)$ and $w(x^{\prime},\xi)=\nabla c(x^{\prime},\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 10 are satisfied. Applying this theorem yields the following guarantee. Let
@@
Next, we set $o(x,\xi)\equiv 0$ and $w(x^{\prime},\xi)=(h^{sm})^{\prime}(c(x^{\prime};\xi))\cdot\nabla c(x^{\prime};\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 10 are satisfied. Applying this theorem provides the following guarantee. Let
@@
At this point, the proof of Theorem 4 is complete.
A.1 Proof of Theorem 10
Our proof of Theorem 10 combines two well-established techniques for proving uniform bounds on the supremum of a random process in the statistical learning theory: chaining, which constructs a sequence of progressively finer approximations to the continuous process, and Sauer-Shelah Lemma, a combinatorial principle using the VC dimension to control the complexity of the function class under consideration [31, Chapter 2]. The chaining argument specifically addresses the continuous parameter $z_{2}$, and the complexity of its increments is then controlled using the VC dimension technique applied to the function class specified by the parameters $z_{1},t$. To deal with measurability issues, we use the calculus of outer integral [31, Chapter 1.2]. A crucial part of our proof demonstrates that our process becomes measurable after symmetrization, building on ideas in the monograph [31, Chapter 2.3].
@@ A.1.1 Chaining Argument
Let $\epsilon_{k}=r2^{-k}$ for every $k\geq 0$. An $\epsilon_{k}$-cover of $\mathcal{Z}_{2}$, denoted by $\mathcal{Z}_{2,\epsilon_{k}}$, enables the approximation of every point $z_{2}\in\mathcal{Z}_{2}$ by a point $\pi_{k}(z_{2})\in\mathcal{Z}_{2,\epsilon_{k}}$ such that $\left\|{z_{2}-\pi_{k}(z_{2})}\right\|\leq\epsilon_{k}$. Specifically, we take the initial cover $\mathcal{Z}_{2,\epsilon_{0}}=\{z_{2,0}\}$ to consist solely of the center $z_{2,0}$ of the ball $\mathcal{Z}_{2}$. By employing a standard volume-type argument, we can further select the cover $\mathcal{Z}_{2,\epsilon_{k}}$ such that $\log|\mathcal{Z}_{2,\epsilon_{k}}|\leq p\log(2+r/\epsilon_{k})$ for every $k\geq 1$ [44, Corollary 4.2.13].
Using these coverings, we can decompose $R_{m}(z_{1},z_{2},t)$ through a chaining argument [45]:
@@
The proof of Lemma 3 is deferred to Section A.1.2.
@@
for each $k\geq 1$. Below we use $C$ to denote the universal constant that appears in Lemma 3.
First, by Lemma 3, with an inner probability at least $21-\delta/$ 2:
@@
Next, the adjustment of $\delta$ into $\delta_{k}$ allows the use of a union bound to extend the probability guarantees from Lemma 3 from individual pairs to all pairs within the cover sets $\mathcal{Z}_{2,\epsilon_{k}}$ and $\mathcal{Z}_{2,\epsilon_{k-1}}$. Specifically, for each $k\geq 1$, with an inner probability at least $1-\delta\cdot 2^{-(k+1)}$, there is the bound:
@@
holds with probability at least $1-\delta$. The proof of Theorem 10 is then complete.
A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
We give a unifying proof for the two statements in Lemma 3. To simplify the notation, we define:
@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
We will first demonstrate how Lemma 5 implies Lemma 3 and defer the proof of Lemma 5 to the end. Introduce two notation:
@@
Note the similarities between the expressions in equation (A.6) and in equation (A.7). Because $u_{z_{2},z_{2}^{\prime}}(\xi)$ and $u_{z_{2,0}}(\xi)$ are subexponential random vectors, the conditions of Lemma 3 are met. More precisely, $u_{z_{2},z_{2}^{\prime}}(\xi)$ is $\sigma\left\|{z_{2}-z_{2}^{\prime}}\right\|$ subexponential for every pair of $z_{2},z_{2}^{\prime}\in\mathcal{Z}_{2}$, and $u_{z_{2,0}}(\xi)$ is $\sigma$ subexponential. This allows us to leverage Lemma 5 to conclude that Lemma 3 follows.
In the remainder, we prove Lemma 5. Our proof exploits the fact that the VC dimension of the set family $\mathcal{W}$ controls the Rademacher complexity associated with the indicators $\psi_{z_{1},t}(\xi)$. Additionally, we utilize calculus rules for outer probability measures [31, Section 1.2-5] to address measurability issues. Recall the notion of outer integral [31, Section 1.2].
Definition 5.
@@
Outer integrals exhibit monotone property: if $T_{1}(\xi)\leq T_{2}(\xi)$ for every $\xi$, then $\mathbb{E}^{*}[T_{1}]\leq\mathbb{E}^{*}[T_{2}]$. This is essentially all we need, as the remaining properties of outer integral will be reference from the monograph [31]. To ensure clarity when discussing the calculus rules of the outer integral, we will explicitly reference the relevant chapters in the monograph [31].
Proof of Lemma 5.
We define the moment-generating function, using the notion of outer integral (see Definition 5):
@@
Let $V_{1/2}$ denote a $1/2$ cover of unit ball in $\mathbb{R}^{d}$ with $\log|V_{1/2}|\leq d\log 6$. Following a standard argument (see, e.g., [44, Exercise 4.2.2]), we know that for every $\bar{v}\in\mathbb{R}^{d}$:
@@
Furthermore, we introduce i.i.d. Rademacher random variables $\{\varepsilon_{i}\}_{i=1}^{m}$ with $\mathbb{P}(\varepsilon_{i}=\pm 1)=1/2$, independent of the data samples $\{\xi_{i}\}_{i=1}^{m}$. To simplify our notation, we use $\varepsilon_{1:m}$ to denote the vector $(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{m})$ and $\xi_{1:m}$ to denote the vector $(\xi_{1},\xi_{2},\ldots,\xi_{m})$.
Applying symmetrization techniques from empirical process theory for the outer integral [31, Lemma 2.3.1], we deduce an upper bound for $U(\tau)$:
$\begin{split}U(\tau)\leq\mathbb{E}^{*}\left[\exp\left(4\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$ (A.10)
where $\mathbb{E}^{*}$ is understood as the outer integral on the product probability space involving the data samples $\xi_{1:m}$ and the independent Rademacher random variables $\varepsilon_{1:m}$. To be more formal, we follow [31, Chapter 2.3]. Recall that a single random variable $\xi$ can be defined on the probability space $(\Xi,\mathscr{G},\mathbb{P})$. Thus, we can define our data samples $\xi_{1:m}$ on the product probability space $(\Xi^{\otimes m},\sigma(\mathscr{G}^{\otimes m}),\mathbb{P}^{\otimes m})$, where $\Xi^{\otimes m}$ represents the $m$-fold Cartesian product of $\Xi$, $\sigma(\mathscr{G}^{\otimes m})$ represents the $\sigma$-field generated by the $m$-fold product of $\mathscr{G}$, and $\mathbb{P}^{\otimes m}$ is the $m$-fold product of $\mathbb{P}$, where $\xi_{1},\xi_{2},\ldots,\xi_{m}$ corresponds to the coordinate projections. Then, we introduce Rademacher random variables $\varepsilon_{1:m}=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{m})$ defined on another probability space $(\Xi^{\prime},\mathscr{H},\mathbb{P}^{\prime})$. We finally combine these to define the final product probability space of interest:
@@
In this product probability space, the sample $\xi_{1:m}$ corresponds to the first $m$ coordinates projections and the additional Rademacher random variables $\varepsilon_{1:m}$ depending only on the $(m+1)$st coordinate. The outer integral $\mathbb{E}^{*}$ in equation (A.10) then refers to the outer integral for this probability space.
Lemma 6 is the critical component that resolves the measurability issue. From the empirical process theory, we know that measurability of certain forms is needed at this point, as Fubini’s theorem is not valid for outer integrals [31, Chapter 2.3]. Lemma 6 follows the established ideas to resolve the measurability issues, see, e.g., discussions in the monograph on the measurable class [31, Definition 2.3.3]. The proof of Lemma 6 is deferred to Section A.1.4.
Lemma 6.
@@
$(\xi_{1:m},\varepsilon_{1:m})\mapsto\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})$
@@
Lemma 6 provides the justification for replacing the outer integral $\mathbb{E}^{*}$ on the RHS of equation (A.10) with the standard integral $\mathbb{E}$. Indeed, equation (A.10) now leads to
$\begin{split}U(\tau)\leq\mathbb{E}\left[\exp\left(4\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$ (A.12)
@@
$\begin{split}V(\tau,\xi_{1:m})=\mathbb{E}_{\varepsilon_{1:m}}\left[\exp\left(\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}\max_{v\in V_{1/2}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\psi_{z_{1},t}(\xi_{i})u_{v}(\xi_{i})\right)\right]\end{split}$
where the symbol $\mathbb{E}_{\varepsilon_{1:m}}$ means the standard expectation taken solely over the Rademacher variables $\varepsilon_{1:m}$. As the Fubini’s theorem holds for the standard expectation, under this notation, the inequality (A.12) then becomes:
@@
For a reference of Sauer-Shelah Lemma, see, e.g., [46, 47] or [44, Theorem 8.3.16].